Find three consecutive integers such that the square of the first plus the product of the other two is 67.
Answer by Alan3354(69443)

Find three consecutive integers such that the square of the first plus the product of the other two is 67.
------------------
Smallest number = x
x^2 + (x+1)*(x+2) = 67
2x^2 + 3x + 2 = 67
2x^2 + 3x - 65 = 0
(x - 5)*(2x + 13) = 0
x = -6.5 Ignore this one, it's not an integer
x = 5